According to Wikipedia, the best-selling album in 2005 was Mariah Carey’s The Emancipation of Mimi. The best-selling album in 2004 was the album Confession by Usher. Confession sold 0.94 million copies less than twice the amount The Emancipation of Mimi sold. The sum of these two albums was 13.97 million copies. Determine the sales of both albums
Answer by stanbon(75887)

The best-selling album in 2004 was the album Confession by Usher.
Confession sold 0.94 million copies less than twice the amount The Emancipation of Mimi sold.
The sum of these two albums was 13.97 million copies.
Determine the sales of both albums.
-----
Equations:
c + e = 13.97 million
c = 2e-0.94
----
Substitute for "c" and solve for "e":
2e-0.94 + e = 13.97
3e = 14.91
e = 4.97 million (# of "Emancipation" sold)
---
Since c = 2e-0.94 = 9 million (# of "Confession" sold)
=======================================================
Cheers,
Stan H.